OpenStudy (twizttiez):

Variable y varies directly with x^2, and y = 32 when x = 4. Which graph represents the quadratic variation?

OpenStudy (twizttiez):

@misty1212

OpenStudy (misty1212):

HI!!

OpenStudy (twizttiez):

Hello!

OpenStudy (misty1212):

\[y=kx^2\] we need \(k\)

OpenStudy (twizttiez):

Would k be 32?

OpenStudy (twizttiez):

Or no y is 32

OpenStudy (misty1212):

no dear put \(x=4,y=32\) and solve \[32=k\times 4^2\] for \(k\)

OpenStudy (misty1212):

\[32=16k\\k=?\]

OpenStudy (twizttiez):

Ok soo k=2?

OpenStudy (misty1212):

yes, and the equation is \[y=2x^2\]

OpenStudy (anonymous):

@TwiztTiez After misty finishes helping you, compare this to the previous question and look at the pattern, ok? =]

OpenStudy (twizttiez):

Ok thanks

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (twizttiez):

But wait now that i have k would i do 32=2x4^2??

OpenStudy (twizttiez):

@misty1212

OpenStudy (twizttiez):

??

OpenStudy (twizttiez):

Anyone?

OpenStudy (twizttiez):

@Nnesha

OpenStudy (twizttiez):

Hello?

OpenStudy (twizttiez):

@pitamar is that what i do?

OpenStudy (twizttiez):

Hello!?

OpenStudy (anonymous):

Sorry, I'm back. let me see

OpenStudy (twizttiez):

Ok

OpenStudy (anonymous):

The thing is that you got k out of 32 = k*4^2. that's why you know k=2. The power this gives you is that now you know the equation is: $$ y = 2 \cdot x^2 $$ And you can find the 'y' values for other points. Take a point you see in the graphs, (say x=2 as we did in previous question), plug it in the equation and see what the 'y' value has to be. Now check what graph is matches those (x,y) values.

OpenStudy (twizttiez):

So A?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Now look at the previous question and see that it is pretty much the same and that you understand the idea. ok?

OpenStudy (twizttiez):

Yeah ok

OpenStudy (twizttiez):

I see now thanks!

OpenStudy (anonymous):

Sure np =)